OpenStudy (anonymous):

A sculpture was purchased for $4,000. The value of the sculpture increases at a rate of 4% per year. How much will the sculpture be worth after 10 years?

OpenStudy (anonymous):

5600

OpenStudy (anonymous):

A = 4000(1.04)^10 assuming annual increases A = 4000exp(0.40) assuming continuous increase

OpenStudy (anonymous):

You would multiply 4 by 10 and get 40 % then use a calculator and put in 4000 + 40% and get 5600

OpenStudy (anonymous):

Wait... so which is it :/

OpenStudy (anonymous):

@douglaswinslowcooper what would i do next to find the answer?

OpenStudy (anonymous):

"4% per year" indicates compounding yearly, usually final value = 4000 (1.04)^10 = 5921

OpenStudy (anonymous):

Thank you! So now it is asking:In how many years will the sculpture be worth $10,000? Round to the nearest tenth of a year. Would i then do 10,000(1.04)^t

OpenStudy (anonymous):

10000 = 4000(1.04)^n solve for n take logarithms 10000/4000=2.5 log(2.5) = n log(1.04) n = log(2.5)/log(1.04) = 23.4 years check: : (1.04)^23.4 = 2.5 checks

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

You are very welcome.